Suppose you have $1,000 today and the risk-free rate of interest ( rf) is 2.3%. The equivalent value in one year is closest to:

$965.00

$1030.00

$1,023

$1000.00
Answer #1

Te future value of any amount invested today can be calculated by using the below formulae

V1=P*(1+R/n)^nt where V1 is the value R= interest rate n= number of times compounded t=time

Here P=$1000 R=2.3% N=1 t=1

Now if we assume the principal with grow at the risk free rate, the value in 1 year is given as

V1=P*(1+R)

thus V1=1000*(1+0.023) = $1023

Thus the final value is $1023